OpenStudy (anonymous):

graph the logarithmic function: log sub 8 x

OpenStudy (anonymous):

\[\log_{8} x\]

OpenStudy (whpalmer4):

What's the issue? Don't know how to calculate values of \(\log_8 x\), or something else?

OpenStudy (anonymous):

sorry didn't mean to disappear. lovely windows update rebooted my pc! bleah.

OpenStudy (anonymous):

don't remember how to calculate or graph log functions

OpenStudy (whpalmer4):

Well, you graph it the same way you graph any function — calculate a range of values, and plot them. To calculate \(\log_8\) you probably need to use the change of base formula, unless your calculator will calculate logs to any base you wish. \[\log_b x = \frac{\log_a x}{\log_a b}\] So if your calculator offers common logarithms (just says Log), you can find\[\log_8 x = \frac{\log_{10} x}{\log_{10}8}\]

OpenStudy (whpalmer4):

In case it isn't clear, \[\log_{10}x\]is the common logarithm, often just written \[\log x\]

OpenStudy (anonymous):

my oh-so-lovely instructor simply asked us to graph the function as I gave it: \[\log_{8} x \] with no other givens

OpenStudy (whpalmer4):

Yes, no other givens are required.

OpenStudy (anonymous):

are you ready to shoot her yet? I sswear....I'm half tempted to get a math degree just so I can teach others and NOT confuzzle them!

OpenStudy (whpalmer4):

Sorry, I think this is a reasonable problem.

OpenStudy (anonymous):

then again....getting a math degree is not a bad idea! will give me a fallback position for job opportunities

OpenStudy (anonymous):

i'm just frustrated with me. i should know this stuff.

OpenStudy (whpalmer4):

I agree ;-)

OpenStudy (whpalmer4):

The nifty thing about this problem is that no matter the range of values of x you choose (so long as you don't try to do something like only use negative values, where the log isn't defined), the graph is going to look pretty much the same.

OpenStudy (anonymous):

thankee-sai, phi. i'd forgotten about this website.